# Many (most?) column subset selection criteria

### v1

@@ 2 Relative Volume
We show that relative volume maximization is NP hard (Section 2.1) and admits no PTAS (Section 2.2).
We introduce the concept of relative volume of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$, which we define as
@@
$\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\sqrt{\det(\boldsymbol{C}^{T}\boldsymbol{C})}=\sigma_{1}(\boldsymbol{C})\cdots\sigma_{k}(\boldsymbol{C}).$ (3)
@@
Geometrically, the relative volume is the ratio of the parallelepiped volume to the volume of the cube that contains the parallelepiped. We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal relative volume, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$.
@@ 2.1 Decision problem
Instance: Matrix $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $0<b\leq 1$.
@@
Relative volume maximization is NP hard.
@@
Following the proof of [9, Proof of Theorem 4], the idea is to perform a polynomial-time reduction from the NP-complete problem Exact Cover by 3-Sets (X3C) [17, Section A.3.1] to relative volume maximization.
@@
Each column of $\boldsymbol{A}$ has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. In the reduced instance we set $m=3M$, $k=M$, and $b=1$. Clearly, constructing $\boldsymbol{A}$ requires polynomial time.
According to [9, Proof of Theorem 4], the instance of X3C is true if and only if there exists a submatrix $\boldsymbol{C}_{*}\in\mathbb{R}^{(3M)\times M}$ with orthonormal columns. Lemma 2 implies that this is equivalent to $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Therefore, the instance of X3C is true if and only if $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Thus, we have constructed a polynomial time reduction from an NP complete problem, establishing that relative volume maximization is NP hard.
@@ 2.2 No PTAS for relative volume maximization
This is analogous to the proof of [9, Theorem 5] for volume maximization. We derive an upper bound for $\operatorname{\mathrm{rvol}}(\boldsymbol{C})$ for the reduced instance of X3C when it is false. Lemma 3 implies that the relative volume increases with the removal of columns. Hence, the largest relative volume is determined by finding two columns for a false instance of X3C.
@@
    whose singular values are $2/\sqrt{3}$ and $\sqrt{2/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=2/\sqrt{3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{2/3}$. Lemma 3 implies
@@
    whose singular values are $\sqrt{5/3}$ and $\sqrt{1/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=\sqrt{5/3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{1/3}$. Lemma 3 implies
@@
Considering the overlap of only two sets is sufficient for the following reason. Suppose that $\ell<k$ columns of $\boldsymbol{C}$ overlap, and they are represented by the submatrix $\boldsymbol{C}_{\ell}$. Lemma 3 implies that $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})$. Let $\boldsymbol{C}_{2}$ be any 2-column submatrix of $\boldsymbol{C}_{\ell}$. Again, Lemma 3 implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{2})$. The non-zero elements of $\boldsymbol{C}_{2}$ can only be of the form (6) or (7). Between the two, (6) achieves the largest relative volume $1/\sqrt{2}$, which implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq 1/\sqrt{2}$.
@@ 3 Volume and S-optimality
We show that volume and S-optimality maximization are NP complete (Section 3.1), give a rigorous proof of the PTAS inapproximability threshold for volume maximization (Section 3.2), and show that S-optimality maximization does not admit a PTAS (Section 3.3).
@@
The S-optimality of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ is defined as [32, (3.5)] and [27, (3.10)]
@@
and greedy algorithms are presented for maximizing the S-optimality.
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal S-optimality, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$.
@@ 3.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $0\leq b\leq 1$.
@@
As in (5), we construct the matrix $\boldsymbol{A}\in\mathbb{R}^{3M\times n}$ where each column has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. According to [9, Proof of Theorem 4], the instance of X3C is true if and only if $\boldsymbol{A}$ has a submatrix $\boldsymbol{C}\in\mathbb{R}^{(3M)\times M}$ with orthonormal columns. Lemma 1 implies that this is equivalent to $\boldsymbol{C}$ satisfying the equalities $\operatorname{\mathrm{vol}}(\boldsymbol{C})=1$ or $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})=1$. Thus, volume maximization and S-optimality maximization are NP hard.
@@ 3.2 No PTAS for volume maximization
Lemma 1 implies that a matrix with unit-norm columns has maximal volume equal to 1. We derive an inapproximability threshold of $2\sqrt{2}/3$.
@@
    Since the second factor is a scalar and contains the orthogonal project $\boldsymbol{P}\equiv\boldsymbol{I}-\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}^{\dagger}$, we can write it as
@@
    Inserting this into the previous expression for the squared volume gives
@@ 3.3 No PTAS for S-optimality maximization
Lemma 1 implies that a matrix with unit-norm columns has maximal S-optimality of $1$. We derive an inapproximability threshold of $(2\sqrt{2}/3)^{1/k}$.
@@ 4 Two-norm and Schatten p-norms
Special cases include the Frobenius norm $\|\boldsymbol{A}\|_{(2)}=\|\boldsymbol{A}\|_{F}$, and the two-norm $\|\boldsymbol{A}\|_{(\infty)}=\|\boldsymbol{A}\|_{2}$. Here we consider only integer values of $p$.
We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with minimal Schatten $p$-norm, $\max_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$.
@@ 4.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $b>0$.
@@
The proof is the same as that of Theorem 3, except we use Lemma 4 instead.
Remark 8.1 explains the requirement $p>2$: The meaningful bound for $p<2$ is an upper bound instead of a lower bound.
@@ 4.2 No PTAS for two-norm minimization
Lemma 4 implies that a matrix with unit-norm columns has a minimal two-norm equal to 1. We give a rigorous derivation of the inapproximability threshold $2/\sqrt{3}$ from [9, Theorem 5].
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [18, Corollary 8.6.3] follows $\sigma_{1}(\hat{\boldsymbol{C}}_{k})\geq\sigma_{1}(\hat{\boldsymbol{C}}_{k-1})$. Together with the above induction hypothesis this implies
@@ 5.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{A})\geq k$ and unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$, and parameter $b>0$.
@@ 5.2 No PTAS for two-norm minimization of the pseudo-inverse
Lemma 5 implies that the pseudo-inverse of a matrix with unit-norm columns has minimal two-norm of 1. We derive an inapproximability threshold of $\sqrt{3/2}$.
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [18, Corollary 8.6.3] follows $\sigma_{k-1}(\hat{\boldsymbol{C}}_{k-1})\geq\sigma_{k}(\hat{\boldsymbol{C}}_{k})$. Together with the above induction hypothesis this implies
@@ 5.3 No PTAS for Frobenius norm minimization of the pseudo-inverse
Lemma 5 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Frobenius norm of $\sqrt{k}$. We derive an inapproximability threshold of $\sqrt{k+\tfrac{1}{4}}$.
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns, where $\|\boldsymbol{c}\|_{2}=1$. From Lemma 10 follows

### v2

@@ 2 Relative Volume
We introduce the concept of maximal volume, and show that relative volume maximization is NP hard (Section 2.1) and admits no PTAS (Section 2.2).
The relative volume of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$, defined as
@@
$\operatorname{\mathrm{vol}}(\boldsymbol{C})\equiv\sqrt{\det(\boldsymbol{C}^{T}\boldsymbol{C})}=\prod_{j=1}^{k}{\sigma_{j}(\boldsymbol{C})}.$ (3)
@@
Geometrically, the relative volume is the ratio of the parallelepiped volume to the volume of the cube that contains the parallelepiped.
Example
In contrast to the volume, the relative volume has the advantage of being able to detect when a matrix is ill-conditioned. For instance, the matrix
$\boldsymbol{A}=\begin{bmatrix}1/\epsilon&0\\ 0&\epsilon\end{bmatrix}\qquad\text{where}\quad 0<\epsilon\ll 1$
is ill-conditioned with condition number $\kappa_{2}(\boldsymbol{A})=1/\epsilon^{2}$. However, the volume does not detect this, since $\operatorname{\mathrm{vol}}(\boldsymbol{A})=1$, while the relative volume does, $\operatorname{\mathrm{rvol}}(\boldsymbol{A})=\epsilon^{2}=1/\kappa_{2}(\boldsymbol{C})$.
Remark 2.1
The relative volume is inversely proportional to some power of the two-norm condition number. Let $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$. Then
$\left(\frac{1}{\kappa_{2}(\boldsymbol{C})}\right)^{k-1}\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\frac{1}{\kappa_{2}(\boldsymbol{C})}.$
Next, we want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal relative volume, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{rvol}}(\boldsymbol{C})}$.
@@ 2.1 Decision problem
Instance: Matrix $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $0<b\leq 1$.
@@
The decision problem of relative volume maximization is NP hard.
@@
Following the proof of [9, Proof of Theorem 4], we perform a polynomial-time reduction from the NP-complete problem Exact Cover by 3-Sets (X3C) [20, Section A.3.1] to relative volume maximization.
@@
Each column of $\boldsymbol{A}$ has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. In the reduced instance we set $m=3M$, $k=M$, and $b=1$. To construct matrices $\boldsymbol{A}$ with more than $3k$ rows, just append the appropriate number of zero rows. This does not change the singular values of $\boldsymbol{A}$. Clearly, constructing $\boldsymbol{A}$ requires polynomial time.
According to [9, Proof of Theorem 4], the instance of X3C is true if and only if there exists a submatrix $\boldsymbol{C}_{*}\in\mathbb{R}^{(3M)\times M}$ with orthogonal columns. Lemma 2 in Section 8.2 implies that this is equivalent to $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Due to this equivalence and the structure of $\boldsymbol{A}$, determining whether $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$ simply amounts to finding orthogonal columns based on the sparsity pattern of $\boldsymbol{C}_{*}$. Therefore, the instance of X3C is true if and only if $\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{*})=1$. Thus, we have constructed a polynomial time reduction from an NP complete problem, establishing that relative volume maximization is NP hard.
@@ 2.2 No PTAS for relative volume maximization
This is analogous to the proof of [9, Theorem 5] for volume maximization. We derive an upper bound for $\operatorname{\mathrm{rvol}}(\boldsymbol{C})$ for the reduced instance of X3C when it is false. Lemma 3 in Section 8.2 implies that the relative volume increases with the removal of columns. Hence, the largest relative volume is determined by finding two columns for a false instance of X3C.
@@
    whose singular values are $2/\sqrt{3}$ and $\sqrt{2/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=2/\sqrt{3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{2/3}$. Lemma 3 in Section 8.2 implies
@@
    whose singular values are $\sqrt{5/3}$ and $\sqrt{1/3}$. Hence, $\sigma_{1}(\hat{\boldsymbol{C}})=\sqrt{5/3}$ and $\sigma_{2}(\hat{\boldsymbol{C}})=\sqrt{1/3}$. Lemma 3 in Section 8.2 implies
@@
Considering the overlap of only two sets is sufficient for the following reason. Suppose that $\ell<k$ columns of $\boldsymbol{C}$ overlap, and they are represented by the submatrix $\boldsymbol{C}_{\ell}$. Lemma 3 in Section 8.2 implies that $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})$. Let $\boldsymbol{C}_{2}$ be any 2-column submatrix of $\boldsymbol{C}_{\ell}$. Again, Lemma 3 implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{\ell})\leq\operatorname{\mathrm{rvol}}(\boldsymbol{C}_{2})$. The non-zero elements of $\boldsymbol{C}_{2}$ can only be of the form (6) or (7). Between the two, (6) achieves the largest relative volume $1/\sqrt{2}$, which implies $\operatorname{\mathrm{rvol}}(\boldsymbol{C})\leq 1/\sqrt{2}$.
@@ 3 Volume and S-optimality
We show that volume and S-optimality maximization are NP hard (Section 3.1), give a rigorous proof of the PTAS inapproximability threshold for volume maximization (Section 3.2), and show that S-optimality maximization does not admit a PTAS (Section 3.3).
@@
The S-optimality of a matrix $\boldsymbol{C}\in\mathbb{R}^{m\times k}$ with $\operatorname{\mathrm{rank}}(\boldsymbol{C})=k$ is defined as [39, (3.5)] and [31, (3.10)]
@@
and greedy algorithms are presented for maximizing the S-optimality. We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with maximal S-optimality, $\max_{\boldsymbol{C}}{\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})}$.
@@ 3.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $0\leq b\leq 1$.
@@
As in (5), we construct the matrix $\boldsymbol{A}\in\mathbb{R}^{3M\times n}$ where each column has exactly 3 non-zero elements and $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$. According to [9, Proof of Theorem 4], the instance of X3C is true if and only if $\boldsymbol{A}$ has a submatrix $\boldsymbol{C}\in\mathbb{R}^{(3M)\times M}$ with orthonormal columns. Lemma 1 in Section 8.1 implies that this is equivalent to $\boldsymbol{C}$ satisfying the equalities $\operatorname{\mathrm{vol}}(\boldsymbol{C})=1$ or $\operatorname{\mathrm{s_{opt}}}(\boldsymbol{C})=1$. Thus, volume maximization and S-optimality maximization are NP hard.
@@ 3.2 No PTAS for volume maximization
Lemma 1 in Section 8.1 implies that a matrix with unit-norm columns has maximal volume equal to 1. We derive an inapproximability threshold of $2\sqrt{2}/3$.
@@
    Since the second factor is a scalar and contains the orthogonal projector $\boldsymbol{P}\equiv\boldsymbol{I}-\hat{\boldsymbol{C}}_{k}^{T}\hat{\boldsymbol{C}}_{k}^{\dagger}$, we can write the second factor as
@@
    Inserting this into (8) gives
@@ 3.3 No PTAS for S-optimality maximization
Lemma 1 in Section 8.1 implies that a matrix with unit-norm columns has maximal S-optimality of $1$. We derive an inapproximability threshold of $(2\sqrt{2}/3)^{1/k}$.
@@ 4 Two-norm and Schatten p-norms
Special cases include the Frobenius norm $\|\boldsymbol{A}\|_{(2)}=\|\boldsymbol{A}\|_{F}$, and the two-norm $\|\boldsymbol{A}\|_{(\infty)}=\|\boldsymbol{A}\|_{2}$. Here we consider only integer values of $p$. We want to find $k$ columns $\boldsymbol{C}$ of $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with minimal Schatten $p$-norm, $\max_{\boldsymbol{C}}{\|\boldsymbol{C}\|_{(p)}}$.
@@ 4.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
@@
The proof is the same as that of Theorem 3, except we use Lemma 4 in Section 8.3 instead.
Remark 8.1 in Section 8.3 explains the requirement $p>2$: The meaningful bound for $p<2$ is an upper bound instead of a lower bound.
@@ 4.2 No PTAS for two-norm minimization
Lemma 4 in Section 8.3 implies that a matrix with unit-norm columns has a minimal two-norm equal to 1. We give a rigorous derivation of the inapproximability threshold $2/\sqrt{3}$ from [9, Theorem 5].
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [21, Corollary 8.6.3] follows $\sigma_{1}(\hat{\boldsymbol{C}}_{k})\geq\sigma_{1}(\hat{\boldsymbol{C}}_{k-1})$. Together with the above induction hypothesis this implies
@@ 5.1 Decision problems
Given a matrix $\boldsymbol{A}\in{\mathbb{R}}^{m\times n}$ with unit-norm columns $\|\boldsymbol{A}\boldsymbol{e}_{j}\|_{2}=1$, $1\leq j\leq n$; integer $1\leq k\leq\min\{\operatorname{\mathrm{rank}}(\boldsymbol{A}),m/3\}$; and parameter $b>0$.
@@ 5.2 No PTAS for two-norm minimization of the pseudo-inverse
Lemma 5 in Section 8.5 implies that the pseudo-inverse of a matrix with unit-norm columns has minimal two-norm of 1. We derive an inapproximability threshold of $\sqrt{3/2}$.
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns. From singular value interlacing [21, Corollary 8.6.3] follows $\sigma_{k-1}(\hat{\boldsymbol{C}}_{k-1})\geq\sigma_{k}(\hat{\boldsymbol{C}}_{k})$. Together with the above induction hypothesis this implies
@@ 5.3 No PTAS for Frobenius norm minimization of the pseudo-inverse
Lemma 5 in Section 8.4 implies that the pseudo-inverse of a matrix with $k$ unit-norm columns has a minimal Frobenius norm of $\sqrt{k}$. We derive an inapproximability threshold of $\sqrt{k+\tfrac{1}{4}}$.
@@
    Let $\hat{\boldsymbol{C}}_{k}\equiv\begin{bmatrix}\hat{\boldsymbol{C}}_{k-1}&\boldsymbol{c}\end{bmatrix}$ be a matrix with $k$ columns, where $\|\boldsymbol{c}\|_{2}=1$. From Lemma 10 in Section 9 follows